OpenStudy (anonymous):

Look at this expression. 24^x4y _______ 3x^-4y^-1 Which of the following is the simplest form of the expression above? a. 8 b. 8x-16y-1 c. 8x-1y-1 d. 8x8y2

OpenStudy (bibby):

\(\huge \frac{a^m}{a^n}=a^{(m-n)}\)

OpenStudy (anonymous):

okay so, it would be option d? 8x^8y^2 ?

OpenStudy (bibby):

\(\huge \frac{24x^4y}{3x^{-4}y^{-1}}=\frac{24}{3}*x^{(4--4)}*y^{1--1}=8x^8y^2\) good job

OpenStudy (anonymous):

thank you (: